OpenStudy (anonymous):

Evaluate each function at x=5/3 f(x) = x^2 + 3x -2 h(x) = x - 1over 2x + 3

myininaya (myininaya):

So it is telling us to find f(5/3) and h(5/3) so f(x)=x^2+3x-2 Notice this is saying f is the machine and x is the input well f(5/3) is saying to use the f machine and wanting you to use the input is 5/3 So just replace x with 5/3 and evaluate

OpenStudy (triciaal):

just replace x with 5/3 and evaluate f(5/3) = (5/3)^2 + 3(5/3) -2 =